OpenStudy (anonymous):

Nadine baked 2 cakes using a circular cake pan of diameter 12 inches. She plans to make a layered cake with one cake on top of the other with frosting only in between the cakes and on top of the second cake. Layla baked 4 cakes using a circular cake pan of diameter 10 inches. She plans to make a layered cake with each cake on top of the other and frosting only in between the cakes and on top of the fourth cake. The height of frosting in both their cakes is the same. Who requires more frosting?

OpenStudy (anonymous):

hey why hvu blocked me silly

OpenStudy (anonymous):

is it really

OpenStudy (anonymous):

it's nadine

OpenStudy (ganpat):

Given for same thickness, Froasting for 12in cake would be 2 * pi/4 * 12 square.. = 226.08 sq.inch.. Froasting for 10in cake = 4 * pi/4 * 10square = 314 sq.inch.. so u have ur answer..